how do i write a function using f(0)=4 f(5)=19

Found 4 solutions by ikleyn, Boreal, TeachMath, greenestamps:
Answer by ikleyn(52874)

.
There are more than million valid answers to this question.

Therefore, without the exact description of the type of the function, the question MAKES NO SENSE.


Answer by Boreal(15235)

if it is f(x), when x=0 the constant is 4
The x term must be 15, and x=5 and there are 4 of them, so 4*5*some function of x must equal 15
x=15/20 or 3/4.
4(3/4)x=3x
when x=5, 3x=15
Therefore, (3/4)x+4=19
That is consistent with f(0)=4 and 4f(5)=15+4=19

Answer by TeachMath(96)

Points: (0, 4) and (5, 19)
Slope, or m = 3

Linear function: y = 3x + 4

Answer by greenestamps(13208)



You got three very different answers, because you were careless in presenting the problem. When you are asking for help with a problem, it would be polite to be careful to present the problem correctly. You might want to try good grammar and punctuation, even including capitalization where needed.

Presumably, since the topic you chose was "linear equations", you want a linear function; but don't assume the tutors here are going to realize that. The way you asked the question, without specifying a linear function, the tutor who said there are millions of possible answers is right -- there is no end to the list of functions that have f(0)=4 and f(5)=19.

And I can only guess that what you wanted to say was that f(0)=4 and f(5)=19, because, with your sloppy language and no punctuation, what your statement of the problem actually says is f(0)=4f(5)=19. That means f(0) is 19 and 4*f(5) is 19 -- in other words, f(0)= 19 and f(5) = 19/4.

So, even though the topic you chose was linear equations, a good way to ask your question CLEARLY might have been this:

How do I write a linear function using f(0)=4 and f(5)=19?